A balloon is filled with 1.46 g of nitrogen gas (N2) at 0 °C and 1.0 atm (STP). What is the volume of the filled balloon?
1
Expert's answer
2020-11-14T13:57:27-0500

volume of the filled balloon = volume of 1.46 g of nitrogen gas (N2) at 0 °C and 1.0 atm (STP)


Temperature = 0°C

So, T = 273 K

Pressure, P = 1 atm


Mass of nitrogen gas, W = 2.46 g

Molar mass of nitrogen gas, M = 28 g/mol


So, number of moles of nitrogen gas,

n= (2.46/28) mol



From the ideal gas equation

PV = nRT

Or, V = nRT/P

Putting the values, and

R = 0.082 L.atm.mol-1.K-1 we get

So, V = (2.46/28)×0.082×273/1

=1.967 L


volume of 1.46 g of nitrogen gas (N2) at 0 °C and 1.0 atm (STP) = 1.967 L

Hence,

volume of the filled balloon = 1.967 L